OpenStudy (anonymous):

1/2 log2(x+4)-1/2 log2 (x-4)=2

OpenStudy (anonymous):

\[1/2 \log _{2}(x+4)-1/2\log _{2}(x-4)=2\]

OpenStudy (lgbasallote):

perform the power rule on your two logarithms first..what do you have?

OpenStudy (anonymous):

@lgbasallote Is it 2 log2 1/2?

OpenStudy (lgbasallote):

are you asking if that's the final answer? coz im thinking the final answer doesnt have a log

OpenStudy (anonymous):

No. You said perform the power rule... on my first log this is the answer I got. Is this correct? Can you show me the steps please?

OpenStudy (lgbasallote):

power rule states that \(\Large a\log_{b} c = \log_{b} a^c\) so \(\Large \frac{1}{2} \log_{2} (x + 4) = \log_{2} (x+4)^{\frac{1}{2}}\) you understand that right? btw..if you need help on the properties do check this out http://openstudy.com/users/lgbasallote#/updates/4fa45f8fe4b029e9dc34e0b5 might help

OpenStudy (henryblah):

1/2log2(x+4)−1/2log2(x−4)=2 log2(x+4)-log2(x-4)=4 log2(x+4/x-4)=4 x+4/x-4=2^4 Solve as needed.

OpenStudy (mertsj):

\[\log_{2}\frac{\sqrt{x+4}}{\sqrt{x-4}}=2 \]

OpenStudy (mertsj):

Put it in exponential form.

OpenStudy (lgbasallote):

oh it's still log form sorry :P

OpenStudy (mertsj):

\[\frac{\sqrt{x+4}}{\sqrt{x-4}}=4\]

OpenStudy (mertsj):

Square both sides.

OpenStudy (mertsj):

\[\frac{x+4}{x-4}=16\]

OpenStudy (mertsj):

\[x+4=16x-64\]

OpenStudy (mertsj):

Finish it up.

OpenStudy (anonymous):

How did you get 16? I thought it was x+4/x-4 = 2^2 which would be x=4/x-4 = 4.

OpenStudy (mertsj):

\[\frac{\sqrt{x+4}}{\sqrt{x-4}}=2^2\]

OpenStudy (mertsj):

When you square both sides, it eliminates the radicals on the left and squares the 4 on the right. 4 squared is 16

OpenStudy (anonymous):

So then I will get x=4+16x-64 correct?

OpenStudy (anonymous):

@Mertsj and @lgbasallote is my answer 4.5333=x?

OpenStudy (mertsj):

Whatever 68/15 is

OpenStudy (anonymous):

So when I got to 68/15 I am correct at that point as well, correct?

OpenStudy (mertsj):

Yep

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

yw